OpenStudy (anonymous):

answer meh question guyss ;p The distance a body falls from rest varies as the square of the time the body is falling. If a body falls 64 ft. in two seconds, how far will it fall in seven seconds? distance fallen is

OpenStudy (anonymous):

somebody's??

OpenStudy (anonymous):

guyss everybody's here lol anyone know where to start?

OpenStudy (anonymous):

yeah I had 224 but it didn't take it...hmmm I'll msg my teacher! Thanks Sarkar!!

OpenStudy (experimentx):

32*7^2

OpenStudy (mathmagician):

first, you have that h(t)=k*t^2 h(2)=64, therefore 64=4k and k=16. h(7)=16*7^2=784

OpenStudy (anonymous):

ohhh thank you!

OpenStudy (anonymous):

okay,didnt see the first line properly,sorry,@mathmagician-thanks

OpenStudy (anonymous):

thanks guys! really appreciate it!